OpenStudy (a1234):

Find the distance traveled in 14 seconds by an object that is moving with a velocity of v(t) = 11 + 6cos t feet per second. A. 154.8204 B. 156.1704 C. 159.9436 Please show how this problem can be solved without derivatives and integrals, which we have not learned about yet.

OpenStudy (danjs):

Are you supposed to approximate it? although those answers look pretty close and accurate

OpenStudy (a1234):

Yes, it is supposed to be an approximation

OpenStudy (danjs):

using rectangles?

OpenStudy (a1234):

probably, although I'm not sure if we're supposed to use some sort of secant line

OpenStudy (danjs):

the cosine function oscillates between +1 and -1 , so if you think the velocity will be increasing and decreasing from 11+6 to 11- 6

OpenStudy (danjs):

The graph of cos(t) will have zero total area for each succession of pi radians... you see how cosine goes above and below the x axis, so each pi, the total area is zero ?

OpenStudy (a1234):

yes

OpenStudy (danjs):

So on the interval to t=14, it will complete 14/pi = 4.456 times, so up until 4pi maybe consider the velocity just a constant 11

OpenStudy (danjs):

So then estimate the area under cos(t) only for the last little bit, t=4pi = 12.566 to t=14

OpenStudy (a1234):

But then how do I get close to the numbers in the choices?

OpenStudy (danjs):

You can draw rectangles under the cos(t) function , the more the better, and the area of each of those is the distance traveled in that rectangle of time..

OpenStudy (a1234):

Okay, I will try that later. Thank you for the help!

OpenStudy (danjs):

v(t) = 11 + 6cos(t) so the total distance can be aprox, the constant 11 for 14 sec plus, the total approximation using rectangles of that last little bit

OpenStudy (danjs):

t=12.566 , 6*cos(t) = 6 t=14 , 6*cos(t) = 0.8204 Over that little time, the velocity goes from 6 to 0.8204, Using just one rectangle half that height difference - 2.5897 and width t=(14 - 12.566) =1.43 The area here is about 1.43 * 2.5897 = 0.8182 distance adding that to the constant 11 speed for 14 seconds distance = 11 * 14 + 0.8182 = 154.81 total

OpenStudy (danjs):

THe value will get better if you divide the small chunk into more rectangles, and add their areas together, instead of using just 1 rectangle like above

OpenStudy (a1234):

Yes, it is more accurate once it's divided into several rectangles

OpenStudy (danjs):